After the footing in Problem 7.9 was built, the groundwater table rose to a depth of...

  1. After the footing in Problem 7.9 was built, the groundwater table rose to a depth of 0.5 m below the ground surface. Compute the new factor of safety against a bearing capacity failure. Com- pare it with the original design value of 2 and explain why it is different. (use Terzaghi’s formula to solve)

7.9) A bearing wall carries a total unfactored load 220 kN/m. It is to be supported on a 400 mm deep continuous footing. The underlying soils are medium sands with c′ = 0, f′ = 37°, and g = 19.2 kN/m3. The groundwater table is at a great depth. Using ASD, compute the minimum footing width required to maintain a factor of safety of at least 2 against a bearing capacity failure. Express your answer to the nearest 100 mm.
Answer #1

Geven that Toral mtactored load=220 kN/mm. Depth (d )= Hoomm = 001 m. Q=37º, q=1 qol khel mosos i diso conseder Terzaghes meX= 8-tw/2-10w-*)] 33 8= 19.2-404117–10:15.) =)!= 96396-0.939 Enims. trzard Arc = 1962X Dode-o = 742 7768ballon from 0 yhn a0et 5 b=0.700mm) Nowo general water table Re at a depon of hoomm., then po b qa39x0.7-00939 Quet = Znc+ entroidNq +0.5 di bną